A chemist prepares a solution of zinc oxalate ZnC2O4 by measuring out 2.41mg of zinc oxalate into a 150.mL volumetric flask and filling the flask to the mark with water. Calculate the concentration in /molL of the chemist's zinc oxalate solution. Round your answer to 3 significant digits.
Answer #1

Mass of zinc oxalate used = 2.41 mg = 0.00241 g.

Molar mass of zinc oxalate = 153.4 g/mol.

No of moles dissolved = 0.00241/153.4 = 1.57 * 10^-5 moles.

Volume of water used = 150 ml = .15 L.

Molar concentration = 1.57 * 10^-5 / .15 = 1.05 * 10^-4 mol/L